OpenStudy (anonymous):

what's the quotient to 66s^2t-63st/3st

OpenStudy (anonymous):

\[\large \frac{ 66s^{2}t -63st}{ 3st }? \] is this your question

OpenStudy (anonymous):

\[\large \frac{ 66s^{2}t -63st}{ 3st } \implies \frac{ 66s^2t }{ 3st }- \frac{ 63st }{ 3st }\] now you may cancel out terms and simplify :), if you need more help just ask.

OpenStudy (anonymous):

Thank you !